derive the expression for penetrating depth of magnetotelluric signals


1
Expert's answer
2021-03-23T08:09:05-0400

Electric-magnetic phenomena can be represented by the Maxwell's equations. 

"J(t)+\frac{\partial \:}{\partial \:t}\left(D(t\right))"

Solution to Maxwell's equations in simple medium

"E_x = A e^{\u2212kz}+B e^{+kz}"

"H_y = \frac{k}{iω \mu_0}(A e^{−kz}+B e^{+kz})"

The homogenous in 1 D

"E_x = A e^{\u2212kz}"

"k =\sqrt{ iω \mu_0/\rho}"

"H_y = \frac{k}{iω \mu_0}(A e^{−kz})"

At "E_x (z=0)= A \implies E_x(z=\delta) = A/ e"

Therefore, penetrating depth of magnetotelluric signals, "\delta =\sqrt{ \frac{2 \rho}{iω \mu_0}}"

Where "ω= 2 \pi f"